OpenStudy (anonymous):

how many carbon atoms are present in a mole of 12 c

OpenStudy (anonymous):

\[12 mol C*\frac{ 6.02*10^{23}atoms }{ 1molC }=7.2264*10^{24}atoms of Carbon\] Use moles of Carbon and multiply it with Avagadros Number then you get atoms.

OpenStudy (cuanchi):

NO!! the question ask in one mole not in 12 moles, what you show there it is an example or it is the solution of the problem?

OpenStudy (anonymous):

Oh shoot your might be right if she's asking how many atoms there are in 12 grams of carbon then the answer is:\[12gC \times \frac{ 1molC }{ 12.011gC}\times \frac{ 6.022\times10^{23}atoms }{ 1molC }=6.016\times10^{23}atomsC\]